OpenStudy (jenniferjuice):

yoyoyoyoyoyo please help ???????!?!?!?!?! Use the substitution method to solve the following system of equations: 3x + 5y = 3 x + 2y = 0 (1, 1) (6, −3) (6, 1) (1, −3)

OpenStudy (anonymous):

Ok, well first, pick a variable to isolate (x, or y).

OpenStudy (anonymous):

We'll do this for the second equation since it's a bit easier.

OpenStudy (jenniferjuice):

y ? @iambatman

OpenStudy (anonymous):

Yeah sure, that works, so now we want to isolate it as you would a normal equation. Since we're dealing with x+2y=0 Can you get y on the left side on your own?

OpenStudy (jenniferjuice):

wait how do i solve it?

OpenStudy (anonymous):

x+2y=0 subtract x from both sides x+2y=0 -x -x 2y = -x now divide both sides by 2 y= -x/2 Does that make sense?

OpenStudy (anonymous):

Next thing you want to do is take that expression and substitute it into equation 1 (where the y is).

OpenStudy (jenniferjuice):

3x + 5(-x/2) = 3 like that?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Now what's 3x - 5x/2?

OpenStudy (anonymous):

ill brb

OpenStudy (anonymous):

Ok back, did you make any progress? 3x - 5x/2, treat as you would as if it's a regular fraction, what's 3-5/2 =?

OpenStudy (jenniferjuice):

i got -1

OpenStudy (anonymous):

3-5/2?

OpenStudy (anonymous):

Try again, :S

OpenStudy (jenniferjuice):

3-5=-2 -2/2 -1?

OpenStudy (anonymous):

What? it's 5/2..like \[3-\frac{ 5 }{ 2 }\] fraction.

OpenStudy (jenniferjuice):

what? how do i solve that?

OpenStudy (anonymous):

At this point in your math career, I'd assume you would know fractions :\

OpenStudy (jenniferjuice):

3/1 - 5/2 ?

OpenStudy (anonymous):

yes!

OpenStudy (jenniferjuice):

its 1/2

OpenStudy (anonymous):

Yes so 3x-5x/2 = x/2

OpenStudy (anonymous):

\[\frac{ x }{ 2}=3\] solve for x

OpenStudy (jenniferjuice):

x=6?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Now you have x, now we take that expression pick one of the equations 1 or 2, lets pick 2 since it's a bit easier again, x+2y=0 we solved x = 6 so sub that into the equation 6+2y=0 solve for y now.

OpenStudy (jenniferjuice):

y=-3?

OpenStudy (jenniferjuice):

@iambatman

OpenStudy (anonymous):

Yes :D, good job!